### v2

@@ 3 Equilibrium
In this section, I solve the baseline game. I begin by analyzing equilibrium play in the voting subgame. In Section 3.2, I analyze candidates’ incentives and equilibrium behavior, and in Section 3.3 I study the equilibrium of the whole game.
@@ 3.1 Voter Behavior
First I turn to the optimal motivated beliefs of voters. The focus is on situations in which $p_{1}\neq p_{2}$, because otherwise voters choose a winner, but not a winning policy. If $p_{1}\neq p_{2}$ off the equilibrium path, $\mathbf{p}$ is not informative about $\omega$. However, it may be so on the equilibrium path. In a symmetric equilibrium, $p_{1}\neq p_{2}$ is only possible in mixed strategies. Because both candidates hold the same information, assume that, when the state is $\omega$, each chooses $p=1$ when with probability $\rho_{\omega}$ and $p=0$ with probability $1-\rho_{\omega}$. Therefore, if $\rho_{1}=1$ and $\rho_{0}=0$, then both choose the policy that matches the state with probability 1. If $\rho_{0}=\rho_{1}=0$, then candidates never choose $p=1$, irrespective of $\omega$.
@@
To see how a voter optimally chooses her motivated belief $\tilde{\mu}$, take the derivative of $AU(s,\mathbf{p},\tilde{\mu})$ with respect to $\tilde{\mu}$:
$\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}=-\frac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[\beta+\Delta-\kappa(1+\beta)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}$ (4)
If $\kappa=\tilde{\kappa}\equiv\frac{\beta+\Delta}{\beta+1}$, then this is equal to zero, independent of $s$. If $\Delta>1$, then $\tilde{\kappa}>1$, and because $\kappa$ is a belief, it must be true that $\kappa<\tilde{\kappa}$, implying that the sign of (4) is independent of $\kappa$. In particular, if this is the case, then $AU(s,\mathbf{p},\tilde{\mu})$ decreases in $\tilde{\mu}$ if $s>0$, and it increases in $\tilde{\mu}$ if $s<0$. Otherwise, that is, if $\Delta\leq 1$, then whether or not $AU(s,\mathbf{p},\tilde{\mu})$ increases in $\tilde{\mu}$ depends on both $s$ and the belief $\kappa$. Note that for each voter $\kappa$ can be interpreted as a measure of trust in government. $\kappa$ is the distribution of beliefs about the policies that will be enacted. If a voter believes that the government will always enact the policy that matches the true state, then $\kappa=\pi(s,\mathbf{p},\tilde{\mu})$. If, however, there is little trust in government regarding policy choices, then $\kappa\neq\pi(s,\mathbf{p},\tilde{\mu})$, and if trust is very low, we have $\kappa\in\{0,1\}$.
We can now derive the optimally distorted complexity of the issue climate change:
@@
Let $\tilde{\kappa}\equiv\frac{\beta+\Delta}{\beta+1}$.
@@
    If $s=0$ or $\kappa=\tilde{\kappa}$, then generally $\tilde{\mu}^{*}=\mu$.
@@
    If $\Delta>1$, then a voter’s optimal distortion of $\mu$ is (i) $\tilde{\mu}^{*}=\infty$ if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
@@
    If $\Delta=1$, then a voter’s optimal distortion of $\mu$, if $\kappa<1$, is (i) $\tilde{\mu}^{*}=\infty$, if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
@@
    If $\Delta<1$, then a voter’s optimal distortion of $\mu$ is (i) $\tilde{\mu}^{*}=\infty$, if either $s<0$ and $\kappa<\tilde{\kappa}$, or $s>0$ and $\kappa>\tilde{\kappa}$, and (ii) $\tilde{\mu}^{*}=0$, if either $s<0$ and $\kappa>\tilde{\kappa}$, or $s>0$ and $\kappa<\tilde{\kappa}$.
The lemma provides us with an important intermediate result, the optimal distortions of $\mu$ for different voters. We see that beliefs about policy only matter if severe climate change causes only moderate baseline welfare losses. Otherwise, if severe climate change leads to catastrophical welfare losses, then any signal indicating that $\omega=1$ will be interpreted as pure noise and hence completely ignored. To the contrary, any signal indicating $\omega=0$ will be accepted as a perfect indication that climate change is indeed mild, independent of the signal’s strength. Finally, a voter with a signal $s=0$ or a voter with $\kappa=\tilde{\kappa}$ has no incentive to distort her signal’s precision and hence updates her belief about $\omega$ like a perfect Bayesian.
What does this imply for the beliefs about $\omega$ the voters hold? Clearly, if $\tilde{\mu}=0$, then beliefs do not change at all, and $\pi(s,\mathbf{p},0)=q$. The same is true for a voter receiving a signal $s=0$, which is not informative about $\omega$. Moreover, if $\tilde{\mu}=\infty$, then any signal that contains only the slightest bit of information will completely move beliefs to the extremes, and hence $\pi(s,\mathbf{p},\infty)\in\{0,1\}$. Only if $s\neq 0$ and $\kappa=\tilde{\kappa}$ will the belief be a non-constant continuous function of $s$ and it equals $\pi(s,\mathbf{p},\mu)$:
@@
    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}0&\text{if }\left(s<0\wedge\kappa<\tilde{\kappa}\right)\\ \hat{\pi}&\text{if }\left(s>0\wedge\kappa<\tilde{\kappa}\right)\vee\left(s<0\wedge\kappa>\tilde{\kappa}\right)\\ \pi(s,\mathbf{p},\mu)&\text{if }\kappa=\tilde{\kappa}\\ 1&\text{if }\left(s>0\wedge\kappa>\tilde{\kappa}\right)\end{array}\right.$
@@
We now know the beliefs of all voters as functions of their signals and of their policy belief $\kappa$. Corollary 1 shows that if $\Delta>1$, and hence severe climate change has catastrophic consequences, then the equilibrium has a simple structure, and $\kappa$ actually plays no roll. Voters have two different beliefs, $\pi(s,\mathbf{p},\tilde{\mu})=0$ and $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$. What does this imply for voters’ decisions at the ballot? Recall that voters vote sincerely for the alternative that they believe maximizes (1). The expected utility from policy 1 is $u(p=1)=-\pi(s,\mathbf{p},\tilde{\mu})\Delta-(1-\pi(s,\mathbf{p},\tilde{\mu}))$, while from policy 0 she gets $u(p=0)=-\pi(s,\mathbf{p},\tilde{\mu})(\Delta+\beta)$. Hence, the voter cast her ballot for policy 1 iff
$u(p=1)>u(p=0)\Leftrightarrow\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}\equiv\frac{1}{1+\beta}.$ (5)
@@
If the reverse is true, then policy 0 is strictly preferred, and if $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$, then a voter is indifferent. Clearly, $01>\tilde{\pi}>$ 0. Note that by Assumption 1 we have $\tilde{\pi}>q$.
@@
If $\Delta>1$, then voters hold two different beliefs, either $\pi(s,\mathbf{p},\tilde{\mu})=0$ or $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$. A voter with the former belief always votes for policy 0, while the decision at the ballot of the other voter depends on $q$, $\beta$, as well as on the politicians’ strategies $\rho_{0}$ and $\rho_{1}$. If $\hat{\pi}>\tilde{\pi}$, then these voters vote for policy 1, and as a consequence all voters vote informatively, i.e., they vote according to their signals. This means that in each state $\omega$ the policy that matches this state is chosen. If, however, $\hat{\pi}\leq\tilde{\pi}$, then a majority of voters always supports policy 0, implying it wins independent of the state.
Things are slightly different when $\Delta\leq 1$. On the one hand, the equilibrium just described continues to exist also when severe climate change decreases baseline welfare only moderately. In particular, this is the case if $\hat{\pi}\leq\tilde{\pi}$ and $\kappa<\tilde{\kappa}$ for sufficiently many $s$. To the contrary, if $\hat{\pi}\geq\tilde{\pi}$ and $\kappa>\tilde{\kappa}$ for sufficiently many $s$, then policy 1 wins independent of the state. Hence, when severe climate change is not catastrophic, then there might exist an equilibrium in which politicians always campaign on the policy tailored for severe climate change, independent of the true state. It follows that we cannot yet exclude that, when $\Delta\leq 1$, a second inefficient equilibrium may exist.
@@
What about the efficient equilibrium, in which candidates choose the optimal policies given $\omega$? Assume each voters wants to believe her signal and thus has $\kappa<\tilde{\kappa}$ when $s<0$ and $\kappa>\tilde{\kappa}$ when $s>0$. Then every voter votes informatively, and therefore the policy matching the state wins for all $\hat{\pi}\in(0,1)$. Hence, if even severe climate change decreases baseline welfare only moderately, then there always may exist an equilibrium in which voters vote informatively, and hence the correct policy is chosen with probability 1.
@@
Assume $\mathbf{p}\in\left\{\left(0,1\right),\left(1,0\right)\right\}$.
@@
    If $\Delta>1$, then there is a unique subgame equilibrium.
@@
      If $\hat{\pi}\leq\tilde{\pi}$, then $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ for all $s$ and a majority of voters always votes for policy 1.
@@
      If $\hat{\pi}>\tilde{\pi}$, then $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$ and $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=1$ if $s\geq 0$, and a majority always votes for the policy $p_{i}$ that matches the true state $\omega$.
@@
    If $\Delta\leq 1$, then there exist multiple subgame equilibria.
@@
      For any $\hat{\pi}\in(0,1)$, there exists a subgame equilibrium with $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$ and $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=1$ if $s\geq 0$, and a majority of voters always votes for the policy $p_{i}$ that matches the true state $\omega$.
@@
      For $\hat{\pi}\geq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=1$ and policy 1 always wins.
@@
      For $\hat{\pi}\leq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=0$ and policy 0 always wins.
@@
      There is no subgame equilibrium in which the policy $p_{i}$, that does not match the state $\omega$, always wins.
On the positive side, there always exists a subgame equilibrium in which information aggregates, if severe climate change decreases baseline welfare not too much, i.e., if $\Delta$ is small. However, this equilibrium depends on the expectations about the efficiency of political process. If voter believe that the policy suggested by their signal is chosen, then the election leads to efficient results. However, if voters are pessimistic, and believe policy 0 will be chosen no matter what is the true state $\omega$, then this becomes a self-fulfilling prophecy if $\hat{\pi}$ is low. When $\Delta$ becomes larger, and hence severe climate change leads to more dire welfare losses, then the efficient equilibrium disappears when $\hat{\pi}$ is small.
@@ 3.2 Candidates
We now study the platform choices of the candidates as a function of their signals. Of course, when they decide which policies to offer they think ahead which policies are likely to lead to electoral success, and therefore voters’ reactions, in particular the one’s formalized in Proposition 1, matter for candidates’ incentives.
@@
First consider potential mixed-strategy equilibria. Focussing on symmetric equilibria, the probability that each candidate choose policy 1 in state $\omega$ is $\rho_{\omega}$ as before. Note that if $\rho_{1}\neq\rho_{2}$, voters learn from platform choices even if $p_{1}\neq p_{2}$. Choosing such a mixed strategy can only be an equilibrium if it leads to a chance of winning of 50%, because candidates need to be indifferent. The probability to win is 50%, if either both receive exactly half the votes, or if the probability that a majority chooses either candidate is 50%. But note that candidates know the state $\omega$, and thus they can infer the exact distribution of signals voters receive. Therefore, in a mixed-strategy equilibrium we need to have that both receive exactly half the votes. This implies policy 1 is implemented with probability of 50%, independent of the state $\omega$. Hence, we need to have $\kappa=\frac{1}{2}$ for all $s$. If $\tilde{\kappa}<\frac{1}{2}$, any voter with $s>0$ will hold belief $\pi(s,\mathbf{p},\tilde{\mu})=1$ and thus they all vote for 1. But this means that in state $\omega=1$, the candidate offering $p=1$ wins with certainty, and thus this cannot be equilibrium. If $\tilde{\kappa}>\frac{1}{2}$, any voter with $s<0$ will hold belief $\pi(s,\mathbf{p},\tilde{\mu})=0$ and thus they all vote for policy 0. But this means that in state $\omega=0$ the candidate offering $p=0$ wins with certainty, and thus this cannot be equilibrium, either. Finally, if $\tilde{\kappa}=\frac{1}{2}$, any voter will choose $\tilde{\mu}^{*}(s)=\mu$. But then information aggregates by Assumption 2, and thus this cannot be an equilibrium belief, either. Hence, we can conclude that no mixed strategy equilibrium can exist.
What about pure strategy equilibria? In a symmetric equilibrium, we must have $p_{1}=p_{2}$, because candidates hold the same information. Suppose voters observe off-equilibrium that $p_{1}\neq p_{2}$ and both candidates expect voters to vote informatively. Hence, they vote for policy 1 if $s>0$, for policy 0 if $s<0$, and they randomize when $s=0$. Then, in each state $\omega$, the candidate offering the policy that matches the state wins the election. Hence, when choosing which policy platform to offer, candidate $i$ knows that $p_{i}=\omega$ wins the election if $p_{-i}\neq\omega$ and otherwise both win with a probability of 50%. To the contrary, choosing $p_{i}\neq\omega$ wins the election with 50% if also $p_{-i}\neq\omega$ and otherwise loses the election for sure. Hence, $p_{i}=s_{i}$ is the unique optimal strategy. It follows that if sufficiently many voters vote informative, candidates will choose informative policy platforms. To the contrary, if candidates expect a sufficiently large number of voters to not vote informatively when $p_{1}\neq p_{2}$, then there exists a policy $\hat{p}\in\{0,1\}$ that wins the election with certainty, independent of the state $\omega$. Candidates then know they win the election by choosing $p_{i}=\hat{p}$ if $p_{-i}\neq\hat{p}$, and they win with a probability of 50% if $p_{-i}=\hat{p}$. Choosing $p_{i}\neq\hat{p}$ will lose the election for sure if $p_{-i}=\hat{p}$, and hence both candidates are best served by offering $\hat{p}$.
This leads to the next formal result:
@@
There is a unique symmetric equilibrium in the platform choice stage, which is in pure strategies. Candidates choose $p_{i}^{*}=\omega$, if they expect a majority of voters to vote for the policy that matches the state, when $p_{1}\neq p_{2}$. Otherwise, there exists a policy $\hat{p}$ that is expected to win independent of the realization of $\omega$ if $p_{1}\neq p_{2}$, and both candidates choose $p_{i}^{*}=\hat{p}$.

### v5

@@ 3 Equilibrium
In this section, we solve the baseline game. We begin by analyzing equilibrium play in the voting subgame. In Section 3.2, we analyze candidates’ incentives and equilibrium behavior, and in Section 3.3 we study the equilibrium of the whole game.
@@ 3.1 Voter Behavior
First we turn to the optimal motivated beliefs of voters. The focus is on situations in which $p_{1}\neq p_{2}$, because otherwise voters choose a winner, but not a winning policy. If $p_{1}\neq p_{2}$ off the equilibrium path, $\mathbf{p}$ is not informative about $\omega$. However, it may be so on the equilibrium path. In a symmetric equilibrium, $p_{1}\neq p_{2}$ is only possible in mixed strategies. Because both candidates hold the same information, assume that, when the state is $\omega$, each chooses $p=1$ when with probability $\rho_{\omega}$ and $p=0$ with probability $1-\rho_{\omega}$. Therefore, if $\rho_{1}=1$ and $\rho_{0}=0$, then both choose the policy that matches the state with probability 1. If $\rho_{0}=\rho_{1}=0$, then candidates never choose $p=1$, irrespective of $\omega$.
@@
Note that this belief is not properly defined when one of the mixed strategies becomes a pure strategy, i.e., when $\rho_{\omega}\in\{0,1\}$. [^4] The reason is that there cannot be a pure symmetric pure strategy equilibrium in which candidates choose $p_{1}\neq p_{2}$. In a symmetric pure strategy equilibrium, platforms are perfectly informative about $\omega$.
[…]
$\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}=-\frac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\mu s}{\sigma^{2}}}}{\sigma^{2}\left((1-q)(1-\rho_{0}){\rho_{0}}+q(1-\rho_{1})\rho_{1}e^{\frac{2\mu s}{\sigma^{2}}}\right)^{2}}$ (4)
[…]
Lemma 1 states for each voter the optimal distortion $\tilde{\mu}^{*}$ as a function of $s$ and possibly $\kappa$:
@@
Let $\tilde{\kappa}\equiv\frac{b+\Delta}{b+1}$.
@@
    If $s=0$ or $\kappa=\tilde{\kappa}$, then $\tilde{\mu}^{*}=\mu$ for all $\Delta$.
@@
    If $\Delta>1$, then (i) $\tilde{\mu}^{*}$$\infty$ if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
@@
    If $\Delta=1$ and $\kappa<1$, (i) $\tilde{\mu}^{*}$$\infty$, if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
@@
    If $\Delta<1$, then (i) $\tilde{\mu}^{*}$$\infty$, if either [$s<0$ and $\kappa<\tilde{\kappa}$] or [$s>0$ and $\kappa>\tilde{\kappa}$], and (ii) $\tilde{\mu}^{*}=0$, if either [$s<0$ and $\kappa>\tilde{\kappa}$] or [$s>0$ and $\kappa<\tilde{\kappa}$].
The lemma provides us with an important intermediate result. We see that beliefs about policy only matter if severe climate change causes only moderate baseline welfare losses. Otherwise, if severe climate change leads to catastrophical welfare losses, then any signal indicating that $\omega=1$ will be interpreted as pure noise. To the contrary, any signal indicating $\omega=0$ will be accepted as a perfect indication that climate change is indeed mild, independent of the signal’s real strength.
What does this imply for the beliefs about $\omega$ the voters hold? Clearly, if $\tilde{\mu}=0$, then
$\pi(s,\mathbf{p},0)=\frac{q\rho_{1}(1-\rho_{1})}{q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})}.$
This is the same belief a Bayesian voter observing $s=0$ holds. Moreover, if $\tilde{\mu}$$\infty$, then any signal that contains only the slightest bit of information will completely move beliefs to the extremes, and hence $\pi(s,\mathbf{p},\infty)\in\{0,1\}$. Only if $\kappa=\tilde{\kappa}$ will the belief be a non-constant continuous function of $s$ and it equals $\pi(s,\mathbf{p},\mu)$:
@@
    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}\pi(s,\mathbf{p},\mu)&\text{if }\kappa=1\\ \hat{\pi}&\text{if }s\geq 0\wedge\kappa<1\\ 0&\text{if }s<0\wedge\kappa<1\end{array}\right$.
  3. (c)
    If $\Delta<1$,
    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}\pi(s,\mathbf{p},\mu)&\text{if }\kappa=\tilde{\kappa}\lor s=0\\ 0&\text{if }s<0\wedge\kappa<\tilde{\kappa}\\ \hat{\pi}&\text{if }\left(s>0\wedge\kappa<\tilde{\kappa}\right)\vee\left(s<0\wedge\kappa>\tilde{\kappa}\right)\\ 1&\text{if }s>0\wedge\kappa>\tilde{\kappa}\end{array}\right$.
@@
We now know the beliefs of all voters as functions of their signals and of their policy belief $\kappa$. Corollary 1 shows that if $\Delta>1$, and hence severe climate change has catastrophic consequences, then the equilibrium has a simple structure, and $\kappa$ actually plays no roll. Voters have two different beliefs, $\pi=0$ and $\pi=\hat{\pi}$. What does this imply for voters’ decisions at the ballot? Recall that voters vote sincerely for the alternative that they believe maximizes (1). The expected utility from policy 1 is $u(p=1)=-\pi\Delta-(1-\pi)$, while from policy 0 she gets $u(p=0)=-\pi(\Delta+b)$. Hence, the voter cast her ballot for policy 1 iff
$u(p=1)>u(p=0)\Leftrightarrow\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}\equiv\frac{1}{1+b}.$ (5)
@@
If the reverse is true, then policy 0 is strictly preferred, and if $\pi=\tilde{\pi}$, then a voter is indifferent. Clearly, $1>\tilde{\pi}>0$. Note that by Assumption 1 we have $\tilde{\pi}>q$.
@@
If $\Delta>1$, then voters hold two different beliefs, either $\pi=0$ or $\pi=\hat{\pi}$. A voter with the former belief always votes for policy 0, while the decision at the ballot of the other voter depends on $q$, $b$, as well as on the politicians’ strategies $\rho_{0}$ and $\rho_{1}$. If $\hat{\pi}>\tilde{\pi}$, then these voters vote for policy 1, and as a consequence all voters vote informatively, i.e., they vote according to their signals. This means that in each state $\omega$ the policy that matches this state is chosen. If, however, $\hat{\pi}\leq\tilde{\pi}$, then a majority of voters always supports policy 0, implying it wins independent of the state. If $\Delta=1$, beliefs are the same except when $\kappa=1$, in which case a voter holds a Bayesian belief about $\omega$ and votes accordingly.
Things are slightly different when $\Delta<1$, as now $\kappa$ starts to matter. As before, a situation in which all voters hold a belief $\kappa<\tilde{\kappa}$, leading to the above discussed situation of all voters voting for policy 0 if faced with a choice, continues to exist. On the other hand, if $\hat{\pi}\geq\tilde{\pi}$ and $\kappa>\tilde{\kappa}$ for sufficiently many $s$, then policy 1 wins independent of the state.
@@
What about an efficient outcome, in which voters collectively choose the policy that matches the state $\omega$? Assume each voters wants to believe her signal and thus has $\kappa<\tilde{\kappa}$ when $s<0$ and $\kappa>\tilde{\kappa}$ when $s>0$. Then every voter votes informatively, and therefore the policy matching the state wins for all $\hat{\pi}\in(0,1)$. Hence, if even severe climate change decreases baseline welfare only moderately, then there may exist an equilibrium in which voters vote informatively, and hence the correct policy is chosen with probability 1.
@@
Assume $p_{1}\neq p_{2}$.
@@
    If $\Delta\geq 1$, then there is a unique equilibrium.
@@
      If $\hat{\pi}\leq\tilde{\pi}$, then $\kappa^{*}=0$ for all $s$ and a majority of voters votes for policy 0.
@@
      If $\hat{\pi}>\tilde{\pi}$, then $\kappa^{*}=0$ if $s<0$ and $\kappa^{*}=\hat{\pi}$ if $s\geq 0$, and a majority votes for $p=\omega$.
@@
    If $\Delta<1$, then there exist multiple equilibria.
@@
      For any $\hat{\pi}\in(0,1)$, there exists a subgame equilibrium with $\kappa^{*}=0$ if $s<0$ and $\kappa^{*}=1$ if $s>0$, and a majority of voters always votes for policy $p=\omega$.
@@
      For $\hat{\pi}\geq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa^{*}=1$ and policy 1 always wins.
@@
      For $\hat{\pi}\leq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa^{*}=0$ and policy 0 always wins.
@@
      There is no subgame equilibrium in which policy $p\neq\omega$ always wins.
There always exists a subgame equilibrium in which information aggregates, if severe climate change decreases baseline welfare not too much, i.e., if $\Delta$ is small. However, this equilibrium depends on the expectations about the efficiency of the political process. If voter believe that the policy suggested by their signal is chosen, then the election leads to efficient results. But if voters are pessimistic, and believe policy 0 will be chosen no matter what is the true state $\omega$, then this becomes a self-fulfilling prophecy if $\hat{\pi}$ is low. When $\Delta$ becomes larger, and hence severe climate change leads to more dire welfare losses, then the efficient equilibrium disappears when $\hat{\pi}$ is small.
@@ 3.2 Candidates
We now study the platform choices of the candidates as a function of their signals. Of course, when they decide which policies to offer, they reason forward thinking about which policy is more likely to lead to electoral success. Therefore voters’ reactions to policy choices, in particular the one’s formalized in Proposition 1, matter for candidates’ incentives.
@@
First consider potential mixed-strategy equilibria. Focussing on symmetric equilibria, the probability that each candidate choose policy 1 in state $\omega$ is $\rho_{\omega}$ as before. If $\rho_{1}\neq\rho_{2}$, voters learn from platform choices even if $p_{1}\neq p_{2}$. Choosing such a mixed strategy can only be an equilibrium if it leads to a chance of winning of 50%, because candidates need to be indifferent. The probability to win is 50%, if either both receive exactly half the votes, or if the probability that a majority chooses either candidate is 50%. But note that candidates know the state $\omega$, and thus they can infer the exact distribution of signals voters receive. Therefore, in a mixed-strategy equilibrium we need to have that both receive exactly half the votes. This implies policy 1 is implemented with probability of 50%, independent of the state $\omega$. Hence, we need to have $\kappa=\frac{1}{2}$ for all $s$. If $\tilde{\kappa}<\frac{1}{2}$, any voter with $s>0$ will hold belief $\pi=1$ and thus they all vote for 1. But this means that in state $\omega=1$, the candidate offering $p=1$ wins with certainty, and thus this cannot be equilibrium. If $\tilde{\kappa}>\frac{1}{2}$, any voter with $s<0$ will hold belief $\pi=0$ and thus they all vote for policy 0. But this means that in state $\omega=0$ the candidate offering $p=0$ wins with certainty, and thus this cannot be equilibrium, either. Finally, if $\tilde{\kappa}=\frac{1}{2}$, any voter will choose $\tilde{\mu}^{*}(s)=\mu$. But then information aggregates by Assumption 2, and thus also this cannot be an equilibrium belief. Hence, we can conclude that no mixed strategy equilibrium can exist.
[…]
This leads to the next result:
@@
For given $\tilde{\mu}^{*}(s)$ for all $s$, there is a unique symmetric equilibrium in the platform choice stage, which is in pure strategies. Candidates choose $p_{i}^{*}=\omega$, if they expect a majority of voters to vote for the policy that matches the state, when $p_{1}\neq p_{2}$. Otherwise, there exists a policy $\hat{p}$ that is expected to win independent of the realization of $\omega$ if $p_{1}\neq p_{2}$, and both candidates choose $p_{i}^{*}=\hat{p}$.